### v1

Abstract
Utilizing the framework of free probability, we analyze the spectral and operator statistics of the Rosenzweig–Porter random matrix ensembles, which exhibit a rich phase structure encompassing ergodic, fractal, and localized regimes. Leveraging subordination formulae, we develop a perturbative scheme that yields semi-analytic expressions for the density of states up to second order in system size, in good agreement with numerical results. We compute higher-point correlation functions in the ergodic regime using both numerical and suitable analytic approximations. Our analysis of operator statistics for various spin operators across these regimes reveals close agreement with free probability predictions in the ergodic phase, in contrast to persistent deviations observed in the fractal and localized phases, even at late times. Notably, the fractal phase exhibits partial features of asymptotic freeness while retaining memory of the initial spectrum, highlighting the importance of non-localized eigenstates for the emergence of free probability behavior. Employing distance measures and statistical tools such as the $\chi^{2}$ statistic, Kullback–Leibler divergence, and Kolmogorov–Smirnov hypothesis testing, we define a characteristic time scale – the free time – that marks the onset of the validity of free probability predictions for operator spectral statistics in the ergodic phase. Remarkably, our findings demonstrate consistency across these different approaches.
1 Introduction
While chaos in the classical domain is a well-defined phenomenon, particularly through the understanding of initial perturbations in phase space trajectories, its extension to the quantum domain presents formidable challenges. This is largely due to the breakdown of the notion of well-defined trajectories in phase space beyond the semiclassical regime, necessitating an entirely different formulation. As such, quantum chaos has traditionally been studied through correlations among the eigenvalues of the Hamiltonian, an approach pioneered by the seminal works of Wigner, Dyson, and Mehta Wigner1 ; Dyson1962a ; mehta1960statistical . These efforts sparked a wave of research, laying the foundation for studying chaotic quantum systems through statistical measures berry77 ; berry85 ; Guhr:1997ve ; Oganesyan:2007wpd ; Atas2013distribution ; Brezin1997 ; Cotler:2016fpe ; Gharibyan:2018jrp .
In addition, recent advancements have shed light on quantum chaos through the study of operator growth. Here, through the Heisenberg time evolution, localized quantum information encoded in simple operators spreads throughout a system, known as information scrambling Roberts:2018mnp . In particular, the time evolution of operators and their correlation functions Hosur:2015ylk ; Cotler:2017jue , often described in terms of diagnostic tools, such as out-of-time-order correlators (OTOCs) LarkinOTOC , and Krylov complexity Parker:2018yvk , have proven useful tools to diagnose quantum chaotic behavior. OTOCs, for instance, measure the sensitivity of a quantum system to initial perturbations, mirroring the classical intuition of phase space trajectories. Krylov complexity, on the other hand, traces the operator through a specific basis and recasts the problem into a particle-hopping problem in a one-dimensional chain Nandy:2024htc . Apart from quantum many-body systems, these studies have been forefront of the holographic understanding of black holes, which are known to be maximally chaotic systems in nature Sekino:2008he ; Maldacena:2015waa .
Since the pioneering work by Wigner, Dyson, and Mehta, random matrix theory (RMT) has emerged as a powerful tool to model chaotic dynamics across various complex physical systems, from nuclear spectra to many-body systems mehta1991random ; haake1991quantum ; PhysRevLett.52.1 . In the context of quantum chaos, RMT has provided deep insights into the statistical properties of energy levels and their correlation, underpinning phenomena such as spectral rigidity and the universal signatures of chaos in closed and open quantum systems. Notably, random matrix ensembles serve as a bridge to understanding how chaotic dynamics give rise to thermalization—an essential principle encapsulated by the Eigenstate Thermalization Hypothesis (ETH) Srednicki:1994mfb ; Deutch_ETH ; DAlessio:2015qtq . In particular, it becomes crucial to extend RMT beyond traditional ensembles, accommodating systems that exhibit ergodic as well as non-ergodic phases. Such extensions help us to study a versatile range of physical systems, including systems that defy complete thermalization, such as many-body localized states or systems exhibiting slow relaxation dynamics Pal_MBL ; Nandkishore:2014kca .
One model that exemplifies this broadened perspective is the Rosenzweig-Porter (RP) model RPmodel , a prototypical random matrix ensemble with a rich phase structure. The Hamiltonian for the canonical version of this model consists of two terms: a diagonal matrix (whose elements are typically sampled from independent Gaussian distributions) is added to a random matrix drawn from one of the classical Gaussian ensembles. The resulting model is characterized by a single parameter that governs the localization properties of system eigenstates and exhibits three different phases, namely, the ergodic, localized, and fractal phases. In the ergodic phase, eigenstates are extended and uniformly spread across the entire system. Conversely, in the localized phase, eigenstates are confined to a small region, indicating minimal overlap with other states. Additionally, the RP model demonstrates fractal behavior, where eigenstates are neither fully extended nor completely localized, but instead occupy a subset of the available space in a non-uniform manner.
Ever since its proposal as a toy model to explain the behavior of energy levels in atomic spectra RPmodel , the RP model has garnered significant interest as an exemplary framework for studying many-body localization, ergodicity, and quantum chaos BREZIN1996697 ; Altland_RP ; BogomolnyRP ; kravtsov2015random . Experimental findings Zhang:2023rpm have demonstrated its suitability in describing spectral transitions from uncorrelated to correlated spectra Guhr_transition ; Guhr2 ; KunzRP . Various adaptations of the RP model have been explored in the literature, including different Dyson symmetry classes Cadez:2024mjc , generalized versions including circular, log-normal, and Lévy ensembles grp1 ; PoisRP ; Buijsman:2021xbi ; lognormalRP0 ; lognormalRP ; LevyRP ; Safonova:2024ffl ; LevyRP2 , and non-Hermitian generalizations nonHRP . The non-Hermitian case is particularly illuminating, as the presence or absence of the fractal phase depends on whether the diagonal potential is real or complex. Several tools have been employed to investigate this model, such as perturbation theory Altland_RP , supersymmetry techniques Guhr_transition ; Guhr2 ; Truong:2016wcq , entropic quantities Pino_2019 ; KhaymovichKL , adiabatic gauge potentials Cadez:2024mjc , Krylov space methods Bhattacharjee:2024yxj , the replica approach Venturelli:2022hka , and the geometry of Hilbert space Sharipov:2024lah .
In recent years, the theory of free probability has emerged as a powerful framework for re-examining quantum chaos from a novel perspective. Originally developed by Voiculescu voiculescu1986addition ; Voiculescu1991 ; Voiculescu1992 in the context of group isomorphisms and operator algebras, free probability theory was later expanded through its close connections with random matrix theory speicher2019lecturenotesfreeprobability ; mingo ; nica2006lectures ; Rao_Edelman_2006 ; edelmanfree , and began gaining traction in physics literature during the 1990s. Early interest was sparked by the works, among others, Gopakumar:1994iq ; Douglas:1994zu ; Zee:1996qu , who explored its relevance to large $N$ gauge theories and matrix models. More recently, the connection between free probability and generalized versions of the ETH has renewed interest in its role in understanding thermalization in quantum systems Pappalardi:2022aaz ; Pappalardi:2023nsj ; Fritzsch:2024qjn ; Wang:2025mzl ; Vallini:2024bwp ; Fava:2023pac . This resurgence has led to a range of developments, including the application of free probabilistic techniques in quantum gravity, where they offer novel insights into entanglement structure and black hole dynamics, particularly in extensions of the “island” paradigm Penington:2019kki ; Wang:2022ots . Notably, recent works Chen:2024zfj ; Camargo:2025zxr have identified a link between operator statistics and the emergence of asymptotic freeness, while other studies Jindal:2024zcg have suggested freeness as a potential missing link between different avatars of quantum chaos.
In this work, we employ an approach based on free probability theory to study the density of states (DOS) and the operator statistics in the RP model. As the system makes a transition from the ergodic phase to the localized phase, the DOS exhibits a transition from a semicircular distribution to a normal distribution. While evaluating the DOS across the three phases of the RP model presents considerable analytical challenges, we develop a semi-analytic approach leveraging free probability tools, including subordination formulae for the resolvent operator. By employing a perturbative method based on the subordination formula, and using the (inverse of the) system size as the perturbation parameter, we derive an approximate DOS in specific regimes of the RP model.
Although similar techniques have been previously applied to analyze the DOS in various models under constrained settings, our method has the practical advantage that it directly uses the free probability results (which otherwise one usually needs to establish first, e.g. using the replica technique and so on) and is also valid for a general class of Hamiltonians where one of the two terms of the Hamiltonian is Gaussian random matrix. As one of the main results of the paper, we obtain the first two orders corrections to the Gaussian distribution of the DOS of the RP model when the system is in the fractal phase. By numerical analysis, we show that our approximate analytical expression for the DOS is in very good agreement with it.
We further study the statistics of the sum of two deterministic spin operators, one of which is time-evolved by the RP model Hamiltonian. At the initial stage, the eigenvalue distribution of the total operator adheres to the free binomial distribution. However, as time progresses, the eigenvalue distribution exhibits significant deviations from this initial distribution. Specifically, in the ergodic regime, the eigenvalue statistics of the operator sum show a transition from the multinomial distribution to the one predicted by the free probability distribution when the component operators are free. The multinomial distribution (discussed later) is the general case of the binomial distribution, which holds only for spin-half operators. Conversely, in the fractal and localized regimes, the eigenvalue distributions diverge substantially from the predictions obtained from free probability theory. This observation lends credence to the broader assertion that chaos leads to the asymptotic freeness between two operators at late times.
In the ergodic phase, we quantify the evolution of these distributions over time by systematically comparing them at late times, with respect to key governing parameters. To achieve this, we employ three distinct statistical measures: the $\chi^{2}$ function, Kullback–Leibler (KL) divergence– widely recognized as a relative entropy in the context of quantum density matrices– and the Kolmogorov–Smirnov (KS) test. While the $\chi^{2}$ function and the divergence rely on computing a notion of statistical distance between two distributions, the KS test is based on the null hypothesis testing. These characterizations are fundamentally different from each other and provide complementary insights, allowing us to define the notion of free time: the onset of a timescale where the free probability description arises. Our analysis reveals a clear trend: the free time tends to grow progressively as the system transitions from the ergodic phase to the fractal phase. Despite the inherent differences between these statistical measures, they exhibit consistent agreement, underscoring the robustness of our findings.
The present manuscript is structured as follows. Section 2 begins with a brief description of free probability, and we also summarize some known results related to the eigenvalue statistics of the sum of two free operators. We introduce the RP model and its spectral statistics in Sec. 3. It is followed by Sec. 4, where the derivation of the DOS of the RP model using the tools of free probability is presented. In section 5 we compute various $2n$-point correlation functions, which is followed by an in-depth analysis of operator statistics in Sec. 6. The notion of free time is introduced, and its various properties are quantified. Section 7 concludes the paper with a summary and some open-ended questions.
2 Elements of Free Probability theory
Free probability theory is the branch of mathematics that studies non-commutative probability spaces. In this framework, the classical notion of independence between random variables is replaced by the notion of free independence or freeness. In this section, we provide a brief review of free probability theory and how to use it to characterize chaotic dynamics.
Non-commutative probability space:
A non-commutative probability space is a pair $(\mathcal{A},\varphi)$, where $\mathcal{A}$ is a unital *-algebra [^2] A unital algebra $\mathcal{A}$ has a multiplicative identity element $\mathbb{I}$ such that $\mathbb{I}a=a\mathbb{I}=a$ for all $a\in\mathcal{A}$. A *-algebra is an algebra equipped with an involution $a\mapsto a^{*}$ satisfying: $(a^{*})^{*}=a$ (involution), $(a+b)^{*}=a^{*}+b^{*}$ (additivity), $(\lambda a)^{*}=\overline{\lambda}a^{*}$ for $\lambda\in\mathbb{C}$ (conjugate-linearity), and $(ab)^{*}=b^{*}a^{*}$ (reverses order of multiplication). and $\varphi:\mathcal{A}\rightarrow\mathbb{C}$ is a linear functional satisfying the following properties:
  • *-linearity: $\varphi(a^{*})=\overline{\varphi(a)}$ ;
  • Unitality: $\varphi(\mathbb{I})=1$ ;
  • Positivity: $\varphi(aa^{*})\geq 0$ for all $a\in\mathcal{A}$ ;
  • Faithfulness: $\varphi(a)=0\Leftrightarrow a=0$.
Examples
  1. 1.
    Quantum mechanics: A canonical example arises in quantum theory, where $\mathcal{A}$ is the algebra of bounded operators on a Hilbert space $\mathcal{H}$, and $\varphi$ is defined via a quantum state $|\psi\rangle\in\mathcal{H}$ by
    $\varphi(a)=\langle\psi|a|\psi\rangle,\quad\text{for }a\in\mathcal{A}\,.$
    This defines a positive, unital, and *-linear functional on $\mathcal{A}$.
  2. 2.
    Random matrix theory: In this setting, $\mathcal{A}$ consists of random $N\times N$ matrices drawn from an ensemble (e.g., Gaussian Ensembles), and the state is given by the normalized ensemble-averaged trace:
    $\varphi(a)=\frac{1}{N}\mathbb{E}[\operatorname{Tr}(a)]\,.$
    This defines a tracial, positive, and normalized linear functional, which is well-defined in the large-$N$ limit for many standard ensembles of random matrices.
Freeness:
Given a non-commutative probability space $(\mathcal{A},\varphi)$, we say that two elements $a,b\in\mathcal{A}$ are free (or freely independent) if the mixed moments of centered polynomials in $a$ and $b$ vanish whenever the variables alternate and each term is individually centered. More precisely, for all $n\geq 1$ and for any polynomials $P_{1},Q_{1},\dots,P_{n},Q_{n}$ in one variable, the following condition holds when $a$ and $b$ are free  mingo ; speicher2019lecturenotesfreeprobability ; nica2006lectures
$\left\langle(P_{1}(a)-\langle P_{1}(a)\rangle)(Q_{1}(b)-\langle Q_{1}(b)\rangle)\cdots(P_{n}(a)-\langle P_{n}(a)\rangle)(Q_{n}(b)-\langle Q_{n}(b)\rangle)\right\rangle=0\,.$ (1)
where we denote the action of the map by $\varphi(\cdot)=\langle\cdot\rangle$.
In the simplest case where all polynomials are linear, i.e., $P(x)=Q(x)=x$, this definition yields:
$\langle(a-\langle a\rangle)(b-\langle b\rangle)\rangle=0\,,$
$\langle(a-\langle a\rangle)(b-\langle b\rangle)(a-\langle a\rangle)\rangle=0\,,$
$\langle(a-\langle a\rangle)(b-\langle b\rangle)(a-\langle a\rangle)(b-\langle b\rangle)\rangle=0\,,$ (2)
and similar relations hold for higher-order terms. These relations impose specific constraints on joint moments. For example, one finds that:
$\langle ab\rangle=\langle a\rangle\langle b\rangle\,,\quad\langle aba\rangle=\langle a^{2}\rangle\langle b\rangle\,,\quad\langle abab\rangle=\langle a^{2}\rangle\langle b\rangle^{2}+\langle a\rangle^{2}\langle b^{2}\rangle-\langle a\rangle^{2}\langle b\rangle^{2}\,.$ (3)
Thus, freeness serves as a non-commutative analogue of statistical independence, ensuring that mixed moments factor in a structured and predictive way, based solely on the individual moments of $a$ and $b$.
Freeness and mixed cumulants:
The implicit condition for freeness in (1) can be expressed more conveniently in terms of special combinations of the moments, known as the free cumulants. The $n$-th (with $n\geq 1$) free cumulant of a random variable $A$ is defined through the following moment-cumulant relation
$\varphi(A^{n})=\sum_{\pi\in NC(n)}\kappa_{\pi}(A),~{}~{}\text{where}~{}~{}~{}~{}\kappa_{\pi}(A)=\prod_{b\in\pi}\kappa_{|b|}(A)~{},$ (4)
where the sum is over all the non-crossing partitions $\pi$ of the set $\{1,\cdots n\}$, denoted as $NC(n)$, and $b$ denotes the blocks of a certain partition $\pi$. Note that the constraint that the sum is over only non-crossing partitions of the set $S=\{1,\cdots n\}$ is the difference between the formula in (4) and an analogous formula relating moments and cumulants in the classical probability theory. The total number of partitions (both crossing and non-crossing ones) of the set $S$ is given by the Bell numbers, the total number of non-crossing partitions is given by the Catalan numbers defined through the relation $C_{n}=\frac{1}{n+1}\binom{2n}{n}$.
For more than one random variable, the relation in Eq. (4) can be generalized to the following
$\begin{split}\varphi(A_{1}A_{2}\cdots A_{n})&=\sum_{\pi\in NC(n)}\kappa_{\pi}(A_{1},A_{2},\cdots A_{n})~{},\\ \text{where}~{}~{}\kappa_{\pi}(A_{1},A_{2},\cdots A_{n})&=\prod_{b\in\pi}\kappa_{|b|}\Big{(}A_{b(1)},A_{b(2)},\cdots A_{b(n)}\Big{)}~{}.\end{split}$ (5)
Once again, $b=(b(1),b(2),\cdots b(n))$ denotes the blocks of a certain partition $\pi$, and $|b|$ indicates its length.
With this definition of the free cumulants, one can express the condition of freeness between two variables $A$ and $B$ as the statement that all the mixed free cumulants of these two variables are zero mingo ; nica2006lectures , i.e., $\kappa_{n}(C_{1},\cdots C_{n})=0$ for $n\geq 2$, $C_{i}\in\{A,B\}$ for all $i$, and $i$ and $j$ such that $C_{i}=A,C_{j}=B$. One straightforward conclusion of this fact, along with the relation between the moments and free cumulants, is that all non-crossing moments containing free variables get factorized in terms of moments of the individual variables.
Free Additive Convolution:
Having discussed the condition that two mutually free random variables, say $a$ and $b$, should satisfy [^3] Throughout this paper, we assume that these two variables are self-adjoint and have well-defined eigenvalue density., we now describe the procedure for obtaining the free probability prediction for the density of eigenvalues of the sum ($a+b$) of these two operators, which will be used in Sec. 6 to compare with numerically obtained distributions. The eigenvalue distribution of $a+b$ is called the free convolution of the eigenvalue distributions $\rho_{a}(\lambda)$ and $\rho_{b}(\lambda)$, and is usually denoted as $\rho_{a+b}(\lambda)=\rho_{a}(\lambda)\boxplus\rho_{b}(\lambda)$ nica2006lectures .
The first ingredient we need to define is the resolvent operator associated with the operators $a$ and $b$, which is the Cauchy transformation of the DOS of the respective operator. E.g.,
$G_{a}(z)=\int\frac{\rho_{a}(\lambda)}{z-\lambda}~{}\text{d}\lambda\,.$ (6)
Here, the integral is over the support of the distribution $\rho_{a}(\lambda)$ on the real line. One can recover the distribution $\rho_{a}(\lambda)$ from the knowledge of the Cauchy transformation $G_{a}(z)$ by using the so-called Stieltjes inversion formula,
[…]
Next, we notice that the functional inverse of the Cauchy transformation (which we denote here by $\mathcal{B}_{a}(z)$) is singular at $z=0$, due to the presence of a pole. Therefore, one can obtain its regular part as
$G^{-1}(z)\equiv\mathcal{B}(z)~{},~{}~{}~{}R(z)=\mathcal{B}(z)-\frac{1}{z}\,.$ (8)
The quantity $R(z)$ is known as the $R$-transformation of the eigenvalue density. Before proceeding further, we note two points about the $R$-transformation. Firstly, the domain of dependence of $R_{a}(z)$ depends on $\rho_{a}(\lambda)$, and it is well-defined only in a region very close to the origin. Secondly, since the functional inverse $\mathcal{B}(z)$ of the Cauchy transformation can have multiple branches, to obtain a well-defined (physical) eigenvalue density, we impose the condition that $R(0)=0$.
With the above-defined $R$-transformation, to obtain the eigenvalue density of the sum of two operators, we utilize the property that the $R$-transformation of the sum $a+b$ of two self-adjoint operators having compactly supported eigenvalue density is the sum of the $R$-transformations of the individual densities voiculescu1986addition ; nica2006lectures . [^4] For derivation of the additivity property of the $R$-transformation for two random matrices, we refer to Zee:1996qu ; zinn1999adding . Thus,
$R_{a+b}(z)=R_{a}(z)+R_{b}(z)\,.$ (9)
Knowing $R_{a+b}(z)$ from $R_{a}(z)$ and $R_{b}(z)$, one can now follow the backward steps to obtain the density $\rho_{a+b}(\lambda)$; namely, first obtain $\mathcal{B}_{a+b}(z)$ from $R_{a+b}(z)$, subsequently invert it to get the corresponding $G_{a+b}(z)$, which can be finally used in the inversion formula (7) to obtain an expression for the free convolved density $\rho_{a+b}(\lambda)$.
A diagrammatic technique was used in Zee:1996qu to derive the additivity property (9) of the $R$-transformation for the addition of two independent large Gaussian random matrices. By noting that the Cuachy transformation $G(z)$ of the density of eigenvalues (i.e., the Green function), can be written in terms of the one-particle irreducible self-energy $\Sigma(z)$ as a Dyson-Schwinger equation $G(z)=1/(z-\Sigma(z))$, we see that the $R$-transformation is just the self-energy,
$\mathcal{B}\big{(}G(z)\big{)}=\frac{1}{G(z)}+\Sigma(z)~{}~{}\rightarrow~{}~{}\Sigma(z)=R\big{(}G(z)\big{)}\,.$ (10)
2.1 Summary of Known Results on Free Additive Convolution
In this subsection, we present some known results for the free additive convolution of free operators. We refer to Camargo:2025zxr for a detailed derivation of these results.
Sum of spin-1/2 free operators:
Consider two free spin-1/2 operators $a$ and $b$ with spectral distribution
$\rho_{a}(\lambda)=\rho_{b}(\lambda)=\frac{1}{2}\left(\delta(\lambda-1)+\delta(\lambda+1)\right)\,.$ (11)
The spectral distribution of their sum is given by the arcsine distribution
$\rho_{a+b}(\lambda)=\begin{cases}\frac{1}{\pi}\frac{1}{\sqrt{4-\lambda^{2}}}\,,~{}~{}~{}-2\leq\lambda\leq 2\,.\\ 0\,,~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{otherwise}\,.\end{cases}$ (12)
This distribution is expected to arise for operators of the form $A+B(t)$ in chaotic models where the constituent operators have a spectral distribution described by a Bernoulli distribution. This includes a broad class of systems, such as spin chains and SYK-like models Camargo:2025zxr .
Figure 1: The operator eigenvalue distributions predicted by the free probability for $s=1/2$ (left) $s=1$ (middle) and $s=3/2$ (right). The spin $1/2$ and the spin $1$ distribution is given by the arcsine distibution (12) and distribution (14) respectively, while the distrbution for spin $3/2$ can be evaluated using the subordination method.
Sum of spin-1 free operators:
Consider two free spin-1 operators $a$ and $b$ with spectral distribution
$\rho_{a}(\lambda)=\rho_{b}(\lambda)=\frac{1}{3}\left(\delta(\lambda-1)+\delta(\lambda)+\delta(\lambda+1)\right)\,.$ (13)
The spectral distribution of their sum is given by the following distribution Camargo:2025zxr
$\rho_{a+b}(\lambda)=\begin{cases}\bigg{|}\frac{f_{1}^{2}(\lambda)-4-9\lambda^{4}+33\lambda^{2}}{3\sqrt{3}\,\pi\lambda\left(\lambda^{2}-4\right)f_{1}(\lambda)}\bigg{|}~{},~{}~{}~{}-1.9227\leq\lambda\leq 1.9227\,.\\ 0\,,~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{otherwise}\,.\end{cases}$ (14)
where we have defined the function
$f_{1}(\lambda)=\left(8+9\left(3\lambda^{4}-30\lambda^{2}+70\right)\lambda^{2}+27\sqrt{-\lambda^{2}\left(\lambda^{2}-4\right)^{2}\left(9\lambda^{4}-33\lambda^{2}-1\right)}\right)^{1/3}\,.$ (15)
The distribution (14) is expected to appear in chaotic models in which the basic operators or composite operators have spectral density of the form (13).
Sum of spin-$s$ free operators:
For operators with spin larger than unity, it becomes increasingly challenging to obtain analytic formulae for the free additive convolution of the sum of operators, due to the fact that one needs to solve order-$s$ algebraic equations. Fortunately, one can obtain such distributions numerically using subordination function methods; see, for instance speicher2019lecturenotesfreeprobability . Such methods were applied in Camargo:2025zxr for the derivation of the free additive convolution of spin-3/2 and spin-2 operators. Figure 1 shows the respective distribution for spin $1/2$, spin $1$, and spin $3/2$. We review the subordination function method in Appendix B.
2.2 Asymptotic Freeness Produced by Chaotic Dynamics
Given two operators $A$ and $B$ that are not initially free, the time evolution of $B$ under a chaotic Hamiltonian is expected to produce a Heisenberg operator $B(t)=e^{iHt}Be^{-iHt}$ that becomes free from $A$ Jindal:2024zcg ; Chen:2024zfj . This emergence of freeness defines a novel notion of quantum chaos, which can be incorporated into the quantum ergodic hierarchy Camargo:2025zxr .
Freeness arises only in infinite-dimensional Hilbert spaces. When operators $A$ and $B$ are represented as $N\times N$ matrices, freeness can emerge only in the limit $N\to\infty$, a regime associated with asymptotic freeness. For finite $N$, cumulants never vanish completely and retain residual values at late times. However, under chaotic dynamics – i.e., when the Hamiltonian exhibits features akin to random matrix theory – the magnitude of these residual cumulants decreases with increasing $N$, suggesting their vanishing in the $N\to\infty$ limit. In this sense, the finite-$N$ system “knows” about its large-$N$ counterpart.
This connection also manifests in operator statistics. Although true freeness between $A$ and the time-evolved operator $B(t)$ requires infinite $N$, the spectral statistics of $A+B(t)$ at finite $N$ often agree with free probability predictions. This agreement serves as a practical signature – or “smoking gun” – of asymptotic freeness.
3 Rosenzweig–Porter random matrix ensemble
3.1 Hamiltonian and the spectral statistics
The Hamiltonian of the RP model is given by RPmodel
$H_{\mathrm{RP}}=A+\frac{1}{N^{\gamma/2}}B\,.$ (16)
Here, $A$ is an $N\times N$ random diagonal matrix, chosen from independent and identically distributed (iid) elements from a normal distribution with zero mean and variance $\sigma^{2}_{A}=4/N$. On the other hand, $B$ belongs to the Gaussian Orthogonal Ensemble (GOE) with zero mean and variance $\sigma^{2}_{B}=2/N$. However, the Hamiltonian can be chosen from any three symmetry classes of Dyson ensembles Cadez:2024mjc . The Hamiltonian shows rich phase structure across different parametric regimes: the ergodic regime for $0\leq\gamma\leq 1$, non-ergodic and extended fractal phase for $21<\gamma<$ 2, and localized phase for $\gamma>2$. The two transitions, namely the ergodic to fractal phase at $\gamma=1$ and the fractal phase to localized phase at $\gamma=2$, can be found analytically at the large $N$ limit. The DOS transitions from a semicircle distribution supported by the interval $[-2,2]$ for $\gamma=0$ to the normal distribution with zero mean and variance $4/N$ for large $\gamma$.
The distributions of nearest-level spacings and their ratios constitute interesting diagnostics of chaotic Hamiltonians. Denoting the eigenvalues of the $N\times N$ Hamiltonian as $\lambda_{n}$ with $n=1,\cdots,N$, the nearest level spacings are given by the difference between the nearest eigenvalues as $\xi_{k}=\lambda_{k+1}-\lambda_{k}$, with $k=1,\cdots,N-1$. For random matrix theory, such distributions follow universal statistics, given by the Wigner–Dyson distribution Haake:2010fgh :
$p_{\upbeta}(\xi)=\frac{2z_{\upbeta}e^{-(\xi z_{\upbeta})^{2}}(\xi z_{\upbeta})^{\upbeta}}{\Gamma\big{(}\frac{1+\upbeta}{2}\big{)}}\,,~{}~{}~{}z_{\upbeta}=\frac{\Gamma\big{(}1+\frac{\upbeta}{2}\big{)}}{\Gamma\big{(}\frac{1+\upbeta}{2}\big{)}}\,,~{}~{}~{}~{}~{}\upbeta=1,2,4\,,$ (17)
where $\upbeta$ is known as Dyson’s index. The three standard Dyson ensembles follow the statistics corresponding to $\upbeta=1$ (GOE), $\upbeta=2$ (Gaussian unitary ensemble, GUE), and $\upbeta=4$ (Gaussian symplectic ensemble, GSE), respectively. The Poisson statistics is computed separately and corresponds to the $\upbeta=0$ limit in the Gaussian $\upbeta$-ensembles Dumitriu:2002beta ; Bujismanbeta , given by $p_{\upbeta=0}(\xi)=e^{-\xi}$. These ensembles form a one-parameter family of Gaussian ensembles which reduces to the standard Dyson ensemble for $\upbeta=1,2,4$.
[…]
$r_{n}:=\frac{\mathrm{min}(\xi_{n+1},\xi_{n})}{\mathrm{max}(\xi_{n+1},\xi_{n})}\,,~{}~{}~{}~{}~{}~{}\langle r\rangle:=\mathrm{Mean}(r_{n})\,.$ (18)
The quantity $\langle r\rangle$ is known as the $r$-ratio, obtained by the mean values of the $r$-value distribution. For Gaussian $\upbeta$-ensembles, the statistical distribution of $r$-values is known Atas_rvalue :
$p_{\upbeta}(r)=\frac{2}{Z_{\upbeta}}\frac{(r+r^{2})^{\upbeta}}{(1+r+r^{2})^{1+(3/2)\upbeta}}\,\Theta(1-r)\,,~{}~{}~{}~{}\upbeta=1,2,4\,,$ (19)
where $\Theta(1-r)$ is the Heaviside $\Theta$ function restricting the value of $r\in[0,1]$, and the normalization constants are given by $Z_{\upbeta}=4/27\,~{}(\upbeta=1)$, $2\pi/81\sqrt{3}\,~{}(\upbeta=2)$, and $2\pi/729\sqrt{3}\,~{}(\upbeta=4)$ respectively Atas_rvalue . For small $r$, the duistribution $p_{\upbeta}(r)\sim r^{\upbeta}$, which share the similar behavior with the level spacing $(\xi)$ distribution $p_{\upbeta}(\xi)\sim\xi^{\upbeta}$ in the same limit. On the other hand, the large $r$ behavior $p_{\upbeta}(r)\sim r^{-(2+\upbeta)}$ is drastically different compared to the rapid exponential decay in $p_{\upbeta}(\xi)$. The Poisson limit is computed separately and given by $p_{\upbeta=0}(r)=2/(1+r)^{2}$.
[…]
Figure 2 (a) depicts the level spacing distribution, $p(\xi)$, for the RP Hamiltonian across three distinct phases: ergodic ($\gamma=0$), fractal ($\gamma=1.5$), and localized ($\gamma=5$). The system size is set to $N=2^{10}$, with $10^{4}$ independent Hamiltonian realizations. In the ergodic phase, the spacing distribution aligns with the Wigner-Dyson distribution of GOE matrices (Eq. (17) with $\upbeta=1$). Conversely, the localized phase exhibits a Poisson distribution, while the fractal phase presents a mixed distribution characteristic.
Additionally, Fig. 2 (b) illustrates the $r$-ratio for the RP Hamiltonian across various parametric regimes of $\gamma$, encompassing ergodic, fractal, and localized phases. The system sizes chosen are $N=2^{8}$, $2^{10}$, and $2^{12}$, with $5\times 10^{4}$, $10^{4}$, and $10^{2}$ ensembles, respectively. These specific values of $N$ are selected for convenience in later computations involving different spin operators, as described in subsequent sections. The result is consistent with the level spacing distribution in Fig. 2 (a).
For different system sizes, a clear crossover at $\gamma_{c}=2$ is observed, which becomes more prominent after collapsing the data into a single curve. The conclusion drawn is that while the nearest level spacing statistics effectively mark the transition from fractal to localized phases, it fails to adequately capture the transition from ergodic to fractal phases at $\gamma=1$. It is noteworthy that such observations have already been reported in the literature Pino_2019 ; PoisRP ; Buijsman:2023ips ; Bhattacharjee:2024yxj , and efforts have been made to identify the ergodic to fractal transitions using various probes.
(a) Level spacing distribution for different $\gamma$.
(b) Level spacing ratio with $\gamma$.
[…]
3.2 Algebra of observables
In order to fully specify a quantum dynamical system, we need not only the system’s Hamiltonian, but also the associated algebra of observables. A generic quantum dynamical system is defined by a pair $(\mathcal{A},\varphi)$, equipped with a time evolution operator $U(t)$. Given the RP model Hamiltonian $H_{\text{RP}}$, the time evolution operator is given by $U(t)=e^{-itH_{\text{RP}}}$. The Hilbert space of the system is spanned by the eigenvectors of $H_{\text{RP}}$.
We consider the algebra of observables generated by the spin-$s$ Pauli operators $X_{i}^{(s)}$, $Y_{i}^{(s)}$, and $Z_{i}^{(s)}$, defined in Section 6.1, along with the identity operator on the corresponding Hilbert space [^5] Given this initial set of operators, one can construct the associated von Neumann algebra by including limits of sequences of operators that converge in the weak operator topology..
Since we will be interested in the spectral properties of operators and their connection with free probability predictions, we will focus on a map of the form $\varphi(\cdot)=\mathbb{E}\,\text{Tr}(\cdot)/N$, which corresponds to an ensemble average of a infinite temperature thermal state.
4 DOS from free probability
4.1 Cauchy and R-transforms
In this section, we outline a procedure for approximating the DOS of the RP model (16) using tools from free probability theory. In the ergodic phase, corresponding to small values of $\gamma$, the DOS is well approximated by the Wigner semicircle distribution. As $\gamma$ increases and the system enters the fractal phase, the DOS gradually deforms, interpolating between the semicircular shape and a Gaussian profile. In the localized phase, which emerges at large $\gamma$, the DOS is well described by a Gaussian distribution. See Fig. 3 for a schematic illustration of this behavior across the three phases. Here, we are interested in analytically computing the DOS specifically for the fractal phase of the RP model. [^6] For other approaches to the calculation of the DOS of the RP model, we refer to kreynin1995density ; Venturelli:2022hka .
Figure 3: Schematic representation of the DOS in the RP model across its three phases: the ergodic phase (left), characterized by a Wigner semicircle distribution; the fractal phase (center), showing an intermediate form between semicircular and Gaussian; and the localized phase (right), where the DOS is well approximated by a Gaussian distribution.
From the functional inverse of the resolvent of an operator $A$ (i.e., the Cauchy transformation of the density of eigenvalues of the operator $A$), we define the $R$-transformation
$G_{A}^{-1}(z)\equiv\mathcal{B}_{A}(z)=\frac{1}{z}+R_{A}(z)~{}.$ (20)
We note that the expression for the resolvent of the operator $A$ in terms of its DOS, $\rho_{A}(\lambda)$ as
$G_{A}(z)=\int\frac{\rho_{A}(\lambda)}{z-\lambda}~{}\text{d}\lambda~{}.$ (21)
Here, the integral is over the support of the distribution $\rho_{A}(\lambda)$. Note that $G_{A}(z)$ is essentially the Cauchy transformation of $\rho_{A}(\lambda)$.
When the operator $A$ is the Hamiltonian of some quantum mechanical system belonging to some random matrix ensemble, one can relate the resolvent with the time evolution operator as follows. Consider the trace of the ensemble-averaged time evolution operator,
$\mathcal{U}_{H}(t)=\frac{1}{N}\mathbb{E}\big{(}\text{Tr}~{}e^{-iHt}\big{)}~{},$ (22)
which is related to the DOS through a Fourier transformation
$\rho_{H}(\lambda)=\frac{1}{2\pi}\int_{-\infty}^{\infty}\text{d}\lambda~{}e^{-itH}~{}\mathcal{U}_{H}(t)~{}.$ (23)
Now, from the definition of the resolvent in (21), it is easy to see that $G_{H}(z)$ is related to the averaged time evolution operator as BREZIN1996697
$G_{H}(z)=i\int_{0}^{\infty}~{}\text{d}t~{}\mathcal{U}_{H}(t)~{}e^{-itz}~{}.$ (24)
4.2 Addition of two free operators: subordination formulae for the resolvent
Now, for addition of two such operators $A$ and $B$, which are free, from the additivity of the $R$ transformation, we have, for the operator $C=A+B$, the following,
$\mathcal{B}_{C}(z)=R_{C}(z)+\frac{1}{z}\,~{}~{}\rightarrow~{}~{}\mathcal{B}_{A}(z)=\mathcal{B}_{C}(z)-R_{B}(z)\,.$ (25)
Substituting $z\rightarrow G_{C}(z)$ in the argument of the last expression above, we have
$\mathcal{B}_{A}\big{(}G_{C}(z)\big{)}=z-R_{B}\big{(}G_{C}(z)\big{)}\,.$ (26)
Finally, applying $G_{A}$ on both sides, we get a self-consistency equation for $G_{C}(z)$,
$G_{C}(z)=G_{A}\big{(}z-R_{B}\big{(}G_{C}(z)\big{)}\big{)}\,.$ (27)
One can perform an entirely analogous procedure for the operator $B$ and obtain
$G_{C}(z)=G_{B}\big{(}z-R_{A}\big{(}G_{C}(z)\big{)}\big{)}~{}.$ (28)
In the following, we use these formulae to obtain an expression for the DOS for the RP model. The choice between the two formulae for obtaining the density of eigenvalues of $C$ – which can be extracted from $G_{C}(z)$ using the Stieltjes inversion formula – depends on whether the resolvent and the $R$-transformation are more straightforward for the operators $A$ or $B$.
Before proceeding further, we note that a self-consistency equation analogous to the one in (27) for the resolvent of the RP model was obtained in Venturelli:2022hka using the replica trick. The authors also discussed the connection of their formula with the one obtained from the law of addition in free probability. The discussion on the approximate evaluation of the DOS of the RP model below follows closely the discussion in Venturelli:2022hka . However, we have kept the discussion below general enough so that it applies not only for the RP model, but addition of two non-commutative operators, one of which has the Wigner semi-circle as the DOS.
4.3 A perturbative scheme for the resolvent
For most of the choices of the operators $A$ and $B$, the expressions for the corresponding resolvents are sufficiently complicated so that it is difficult to solve either of the equations in eqs. (27) or (28) for $G_{C}(z)$. Therefore, we now proceed as follows to obtain a perturbative expression for the DOS of an operator of the form $C=A+\alpha B$, where $A$ and $B$ are mutually free and the expressions for their DOS are assumed to be known. Here $\alpha$ is a parameter which we assume to be small so that one can set up a perturbation series with respect to this parameter. E.g., for the RP model Hamiltonian, $\alpha=N^{-\gamma/2}$. Using the scaling transformation rule of the $R$-transformation [^7] The $R$-transformation of the density of eigenvalues of an operator $A$ follows the scaling law: $R_{\alpha A}(z)=\alpha R_{A}(\alpha z)$, where $\alpha$ is a constant., and the definition (in eq. (21)), we have, from eq. (27),
$G_{C}(z)$ $=\int\text{d}\lambda~{}\frac{\rho_{A}(\lambda)}{z-\alpha R_{B}\big{(}\alpha G_{C}(z)\big{)}-\lambda}$
$=i\int_{0}^{\infty}~{}\text{d}u~{}\chi_{A}(-u)\exp[-iu\Big{(}z-\alpha R_{B}\big{(}\alpha G_{C}(z)\big{)}\Big{)}\Big{]}~{}.$ (29)
Here we have defined the characteristics function $\psi_{A}(u)$ of the distribution $\rho_{A}(\lambda)$ by a Fourier transformation
$\chi_{A}(u)=\int\text{d}\lambda~{}\rho_{A}(\lambda)~{}e^{-i\lambda u}~{}.$ (30)
Comparing this expression with the inverse of the equation. (23), we see that the characteristic function is nothing but the ensemble-averaged evolution operator generated by $A$, i.e., $\chi_{A}(-t)=\mathcal{U}_{A}(t)$.
Next, expanding the exponential factor containing $R_{B}$ in eq. (4.3) we have the following series expansion for $G_{C}(z)$,
$G_{C}(z)=i\int_{0}^{\infty}~{}\text{d}u~{}\chi_{A}(-u)~{}e^{-iuz}\sum_{n}\frac{(iu\alpha)^{n}}{n!}\big{(}R_{B}\big{(}\alpha G_{C}(z)\big{)}\big{)}^{n}~{}.$ (31)
Writing in terms of the free cumulants ($\kappa_{j}(B)$) of the operator $B$, this has the form
$G_{C}(z)=\sum_{n}\frac{i^{n+1}\alpha^{n}}{n!}\Big{(}\sum_{j=1}^{\infty}\kappa_{j}(B)\big{(}\alpha G_{C}(z)\big{)}^{j}\Big{)}^{n}\int_{0}^{\infty}~{}\text{d}u~{}\chi_{A}(-u)~{}u^{n}~{}e^{-iuz}~{}.$ (32)
So far, this expression for $G_{c}(z)$ is exact. For a general operator $B$, its $R$-transformation can be an arbitrary function of $z$, so that it is difficult to obtain a perturbation series with $\alpha$ as a small parameter. However, one particular case where the above expression becomes easily tractable is when the operator $B$ has Wigner semi-circle as the DOS (which is the case also for the RP model considered in this paper), i.e., $\rho_{B}(\lambda)=\frac{1}{2\pi}\sqrt{4-\lambda^{2}}$. In that case [^8] The Wigner semi-circle is the analogue of the Gaussian distribution for non-commutative probability, i.e., this distribution has only the first free cumulant non-zero. Since $R(z)$ is the generator of the free cumulants, this is equal to $R(z)=z$ for the semi-circle distribution. $R(z)=z$, and the expression in eq. (31) (or (32)) simplifies to
$G_{C}(z)$ $=\sum_{n}\frac{i^{n+1}}{n!}\alpha^{2n}\big{(}G_{C}(z)\big{)}^{n}\int_{0}^{\infty}\text{d}uu^{n}\chi_{A}(-u)e^{-iuz}\,,$
$=\sum_{n}\frac{(-1)^{n}}{n!}\alpha^{2n}\big{(}G_{C}(z)\big{)}^{n}\partial_{z}^{n}G_{A}(z)\,,$ (33)
where to obtain the last expression, we have used an analogous relation to that of the one in (24), valid for the operator $A$.
Evaluating the terms of the above sum recursively, one can obtain the resolvent $G_{C}(z)$ with corrections of different orders of $\alpha$. E.g., at $\mathcal{O}(\alpha^{0})$
$G_{C}^{(0)}(z)=i\int_{0}^{\infty}~{}\text{d}u~{}\chi_{A}(-u)~{}e^{-iuz}~{}=G_{A}(z)~{}.$ (34)
The expressions for $G_{C}(z)$ with first and second order corrections are, respectively,
$G_{C}(z)\approx G_{C}^{(1)}(z)+\mathcal{O}(\alpha^{4})~{},~{}~{}\text{where}~{}~{}~{}~{}G_{C}^{(1)}(z)~{}=G_{C}^{(0)}(z)-\alpha^{2}G_{C}^{(0)}(z)~{}\partial_{z}G_{C}^{(0)}(z)~{}.$ (35)
and $G_{C}(z)\approx G_{C}^{(2)}(z)+\mathcal{O}(\alpha^{6})$ where
$G_{C}^{(2)}(z)$ $=G_{C}^{(0)}(z)-\alpha^{2}G_{C}^{(0)}(z)\partial_{z}G_{C}^{(0)}(z)\,$
$+\frac{\alpha^{4}}{2}\Big{(}G_{C}^{(0)}(z)-\alpha^{2}G_{C}^{(0)}(z)\partial_{z}G_{C}^{(0)}(z)\Big{)}^{2}\partial_{zz}G_{C}^{(0)}(z)\,.$ (36)
Therefore, all orders of correction can be obtained in terms $G_{C}^{(0)}(z)=G_{A}(z)$ (and its various order derivatives with respect to $z$), which in tern can be obtained from the characteristics function of the distribution $\rho_{A}(\lambda)$. Note also that due to the specific scaling law for the $R$-transformation, the $n$-order correction to $G_{C}^{(0)}(z)$ is proportional to $\alpha^{2n}$. The approximate DOS of the operator $C$ can now be obtained from $G_{C}(z)$ using the Stieltjes inversion formula (7),
$\rho_{C}(\lambda)=\frac{1}{\pi}\lim_{\epsilon\rightarrow 0}(\text{Im}~{}G_{C}(\lambda-i\epsilon))~{}.$ (37)
Figure 4: DOS of the RP model: comparison of approximate analytical expression with numerical results. The blue and black-dashed curves respectively represent the analytical DOS with first and second-order corrections. The red curve represents the Gaussian eigenvalue distribution of the matrix $A$. Here we have set $s=1/2,~{}\gamma=1.1,L=10$, and the histogram shows the numerical DOS averaged over 5000 independent realisations of the RP model Hamiltonian.
4.4 Approximate DOS of the RP model
Until now, the approximate expressions for $G_{C}(z)$ we have derived above are valid for any operator $A$ with a well-defined DOS, which is then added with an operator $B$ having a semi-circle density. Now specializing to the RP model considered in this paper, we see that the eigenvalue distribution for the diagonal matrix $A$ is a Gaussian distribution with zero mean and variance $\sigma^{2}=4/N$. Hence, denoting the operator $C$ as the Hamiltonian $H$ of the RP model, we have
$G_{C}^{(0)}(z)=\sqrt{\frac{\pi}{2\sigma^{2}}}\Bigg{[}ie^{-\frac{z^{2}}{2\sigma^{2}}}+\frac{2}{\sqrt{\pi}}D\bigg{(}\frac{z}{\sqrt{2}\sigma}\bigg{)}\Bigg{]}~{},$ (38)
[…]
In Figure 4, we have shown the DOS of the RP model obtained using this procedure along with the one obtained numerically. [^10] In Ref. Truong:2016wcq , an implicit formula for the DOS of a Hamiltonian where a random matrix perturbation ($V$) is added to a random (or a non-random matrix) matrix $H_{0}$ was derived using supersymmetry techniques, including the case of the RP model. The numerical results reported there seem to be consistent with the ones reported here. As can be seen, apart from a small region at the center of the spectrum, the approximate expression has a good agreement with numerical results.
5 Behavior of $2n$-point correlators
5.1 Choice of operators
The $n$-point correlation functions of an operator $\mathcal{O}$ with appropriate normalization that we are interested in are defined as
$F_{2n,(s)}(t)=\langle(\mathcal{O}^{\dagger}(0)\mathcal{O}(t))^{n}\rangle=\frac{1}{(2s+1)^{L}}\mathbb{E}\,\mathrm{Tr}\big{[}(\mathcal{O}^{\dagger}(0)\mathcal{O}(t))^{n}\big{]}\,,~{}~{}n=0,1,2,\cdots\,,$ (39)
where $\mathcal{O}(t)=e^{iHt}\mathcal{O}(0)e^{-iHt}$ is the time-evolved operator and $\langle\cdot\rangle=\mathbb{E}(\mathrm{Tr}(\cdot))/(2s+1)^{L}$. Here, $\mathbb{E}$ refers to the ensemble average, assuming the Hamiltonian $H$ is drawn from a random matrix ensemble. For the RP model, the averaging encompasses the distributions of both operators $A$ and $B$ in (16).
By definition, $F_{0,(s)}(t)=1$ for all time. For example, the $2$-point and $4$-point functions are given for $n=1$ and, $n=2$ respectively
$F_{2,(s)}(t)$ $=\frac{1}{(2s+1)^{L}}\mathbb{E}\,\mathrm{Tr}\big{(}\mathcal{O}^{\dagger}(0)\mathcal{O}(t)\big{)}\,,$ (40)
$F_{4,(s)}(t)$ $=\frac{1}{(2s+1)^{L}}\mathbb{E}\,\mathrm{Tr}\big{(}\mathcal{O}^{\dagger}(0)\mathcal{O}(t)\mathcal{O}^{\dagger}(0)\mathcal{O}(t)\big{)}\,.$ (41)
The $4$-point functions are commonly known as the out-of-time-ordered correlator (OTOC). Since we will be focusing on the unitary evolution of the Hermitian operators, we have $\mathcal{O}^{\dagger}(t)=\mathcal{O}(t)$ for all time.
We define the operators as tensor products of spin operators with the following explicit forms:
$S_{1}$ $\equiv S_{1}(0)=\epsilon\sqrt{\frac{s(s+1)}{3}}\,Z_{1}^{(s)}\otimes I_{2}\otimes\cdots\otimes I_{L}\,,$ (42)
$S_{L}$ $\equiv S_{L}(0)=\epsilon\sqrt{\frac{s(s+1)}{3}}\,I_{1}\otimes\cdots\otimes I_{L-1}\otimes Z_{L}^{(s)}\,,$ (43)
where the spin system consists of $L$ lattice sites, with $Z_{k}^{(s)}$ and $I_{k}$ being the Pauli $z$ matrices (of spin $s$) and identity operators acting on site $k$. The parameter $\epsilon$ distinguishes between spin types: $\epsilon=2$ for half-integer spins and $\epsilon=1$ for integer spins.
The operators are normalized using the Frobenius inner product: $\langle P|Q\rangle=\mathrm{Tr}(P^{\dagger}Q)/(2s+1)^{L}$, ensuring the dimensional consistency. Here, $S_{k}$ refers to spin operators corresponding to spin $s$, with dimension $(2s+1)^{L}$. The spin operators adhere to the following fundamental relations, which govern their commutation properties and additive structure acting on site Craps:2019rbj
$[X_{j}^{(s)},Y_{k}^{(s)}]=i\sqrt{\frac{3}{s(s+1)}}\,\delta_{jk}\,Z_{j}^{(s)}\,,$ (44)
$X_{i}^{(s)}X_{i}^{(s)}+Y_{i}^{(s)}Y_{i}^{(s)}+Z_{i}^{(s)}Z_{i}^{(s)}=3\,\mathbb{I}_{N}\,,~{}~{}~{}~{}(i~{}\mathrm{not~{}summed}).$ (45)
where $\mathbb{I}_{N}$ represents the identity matrix of dimension $N=(2s+1)^{L}$, same as the dimension of the RP model Hamiltonian. For the special case of $s=1/2$, the operators reduce to the Kronecker products of Pauli matrices, thereby aligning with the familiar formalism of spin $1/2$.
Using the above definition, we study the behavior of the $n$-point functions
$F_{2n,(s)}^{(\gamma)}(t)=\frac{1}{(2s+1)^{L}}\mathbb{E}\,\mathrm{Tr}\big{[}(S_{1}(0)S_{1}(t))^{n}\big{]}=\langle(S_{1}(0)S_{1}(t))^{n}\rangle\,,~{}~{}n=0,1,2,\cdots\,,$ (46)
throughout the regime $\gamma\in[0,1]$, i.e., in the ergodic phase. Similarly, the evolution can be defined for $S_{L}$. From now on, we concentrate on operators with spin $s=1/2$ only, with different system sizes $L$, and omit the subscript $s$.
5.2 Approximate analytical expressions for the two-point function
In random matrix theory, one can show that the two-point functions of operators are closely related to the spectral form factor (SFF), while the higher-order correlation functions are also related to the higher-order SFFs  Cotler:2017jue . At early times, the SFF is well-approximated by the square of the partition function – corresponding to the disconnected contribution of the correlation function. We expect this relation to hold approximately in the ergodic phase of the RP model, particularly for small values of $\gamma$ [^11] Note that the relation between the two-point function and the SFF that was derived in Cotler:2017jue is only valid for random matrix ensembles having unitary invariance. However, for the RP mode, the rotational invariance is broken, even for small values of $\gamma$. Therefore, the relation in (47) between the two-point function and the disconnected part of the SFF that we propose below is essentially an observation. Following the approach of Cotler:2017jue , we estimate the early-time behavior of the two-point function as:
$F_{2}^{(\gamma)}(t)=\langle S_{1}(0)S_{1}(t)\rangle\approx Z(t)^{2}/N^{2}\,,\qquad Z(t)=N\int\textrm{d}E\,\rho(E)e^{-iEt}\,,\qquad(t\lesssim t_{d})$ (47)
where $t_{d}$ denotes the timescale up to which the approximation remains valid, $N$ is the rank of the Hamiltonian, and $\rho(E)$ is the DOS.
Figure 5: Behavior of 2-point (left), 4-point (middle), and $2n$-point functions of the operator $S_{1}$ (with $s=1/2$) in the RP model. The black dashed line indicates the analytic result (49) for $\gamma\leq 0.5$, which reduces to (50) for $\gamma=0$. We choose $L=8$, and a single realization of the Hamiltonian is taken.
Deep in the ergodic phase of the RP model (with sufficiently small value of the parameter $\gamma$), and sufficiently large $N$, the normalized DOS is well approximated by a Wigner semicircle kravtsov2015random ; guhr1996transitions
$\rho(E)=\frac{2}{\pi R^{2}}\sqrt{R^{2}-E^{2}}\,,$ (48)
where $R$ is the radius of the semicircle distribution. For the RP model, we find that in the regime of large $N$ and small $\gamma$, the radius scales approximately as $R=2/N^{\gamma/2}$, consistent with the structure of the model’s Hamiltonian [^12] For $N=500$, we observed that the Wigner semicircle described well the DOS as long as $\gamma\leq 0.5$..
Substituting Eq. (48) into Eq. (47), we obtain an analytical approximation for the two-point function:
$F_{2}^{(\gamma)}(t)\approx\left(\frac{2J_{1}(Rt)}{Rt}\right)^{2}\,,\qquad R=\frac{2}{N^{\gamma/2}}\,,$ (49)
where $J_{1}$ denotes the Bessel function of the first kind. For $N=500$, we find that this expression accurately captures the early-time behavior of the two-point function for $\gamma\lesssim 0.5$. For $\gamma=0$ (49) reduces to the result derived in Cotler:2017jue :
$F_{2}^{(\gamma=0)}(t)=\frac{J_{1}(2t)^{2}}{t^{2}}\,.$ (50)
Figure 5 (left and middle) shows the numerical results of the $2$-point and $4$-point functions for $L=8$. The analytic expression (49) matches with the numerical results at early times for small values of $\gamma$. These cases are dominated by the off-diagonal part of the GOE Hamiltonian (16). The results obtained with (49) are shown by the black dashed line in the left panel of Fig. 5. As $\gamma$ increases, the decay develops a longer tail. On the other hand, the 4-point functions exhibit clear zeroes and sharp local minima in the time axis for all $\gamma$. The decay is also faster than that of the 2-point functions. To analyze this closely, we study the behavior of $2n$-point functions for $n=1,\cdots,5$ with $\gamma=0$ for a system size of $L=8$. This is illustrated in the rightmost panel of Fig. 5. As observed, for a fixed $\gamma$, higher-point functions exhibit a faster decay.
6 Operator statistics from free probability
In this section, we investigate the statistical properties of operators within the framework of the RP model. Specifically, our focus lies on the eigenvalue statistics of the following operator:
$S(t)=S_{1}(0)+S_{L}(t)=S_{1}(0)+U^{\dagger}(t)S_{L}(0)U(t)\,,~{}~{}~{}U(t)=e^{-iH_{\mathrm{RP}}t}\,,$ (51)
where the temporal evolution is generated by the RP model Hamiltonian, as defined in (16). Here $S_{1}(0)$ and $S_{L}(0)$ represent the spin operators at the initial time $t=0$, defined in (42) and (43). The total sum of the operators $S(t)$ constructed through this formalism, encapsulates the impact of the RP Hamiltonian (16) over time. Our primary objective is to systematically analyze the eigenvalue statistics of $S(t)$ at various time instants $t$. By varying $\gamma$, i.e., the parametric regime of the RP Hamiltonian, distinct phases (ergodic, fractal, or localized) or structures in the operator statistics are anticipated, potentially exhibiting novel characteristics inspired by predictions from free probability theory.
6.1 Choice of the initial operators
In the following analysis, we consider the operators $S_{1}(0)$ and $S_{L}(0)$ to be spin operators associated with spin $s$. This formulation generalizes the construction and enables comparisons against free probability predictions for various spin values, as obtained through analytical methods outlined in Section 2. The operators $S_{1}(0)$ and $S_{L}(0)$ are defined as tensor products of spin operators as in (42)-(43).
Figure 6: The statistics of the eigenvalues of the time-evolved $s=1/2$ operator (51) at the initial time $t=0$ (top, left) and late times $t=50$ in the ergodic (top, right), fractal (bottom, left), and localized (bottom, right) regimes. We choose $L=8$ (i.e., dimension of the Hamiltonian $N=(2s+1)^{L}=256$) with $5000$ ensemble averages. The solid blue lines describe the arsine distribution (12) i.e., Fig. 1 (a).
We analyze three distinct spin values $s=\{1/2,1,3/2\}$, each paired with a corresponding lattice size $L=\{8,6,5\}$. Using the Frobenius norm discussed earlier, the eigenvalues of the operators $S_{1}$ and $S_{L}$ are determined as follows:
Spin $s$ Operators Eigenvalues
$1/2$ $S_{1},S_{L}$$\{1,-1\}$
1 $S_{1},S_{L}$$\{1,0,-1\}$
$3/2$ $S_{1},S_{L}$$\{3,1,-1,-3\}$
Table 1: Eigenvalues of the operators $S_{i}$, with $i=1,L$ for different spin values.
This is a convention that we follow throughout our discussion. More generally, we take any spin $s$ of our choice, with an $(2s+1)^{L-1}$ equal number of degenerate and distinct eigenvalues ranging from $-\epsilon s$ to $\epsilon s$ in steps of $\epsilon$, where $\epsilon=2$ for half-integrer spins and $\epsilon=1$ for integrer spins, as introduced in (42)-(43). It is worth noting that alternative constructions of the higher-spin operators are possible. These constructions may lead to different commutation relations, such as: $[X_{j}^{(s)},Y_{k}^{(s)}]=in_{s}\delta_{jk}\,Z_{j}^{(s)}$, where $n_{s}=1/s$ with the additive relations $X_{i}^{(s)}X_{i}^{(s)}+Y_{i}^{(s)}Y_{i}^{(s)}+Z_{i}^{(s)}Z_{i}^{(s)}=\big{(}\frac{s+1}{s}\big{)}\,I_{N}$, as discussed in Refs. Xu:2019lhc ; Bhattacharjee:2022vlt . Another possible formulation is: $[X_{j}^{(s)},Y_{k}^{(s)}]=i\delta_{jk}\,Z_{j}^{(s)}$ with $X_{i}^{(s)}X_{i}^{(s)}+Y_{i}^{(s)}Y_{i}^{(s)}+Z_{i}^{(s)}Z_{i}^{(s)}=s(s+1)\,I_{N}$, as explored in Ref. Yin:2020oze . Despite these differing constructions, it is straightforward to demonstrate that they are related through appropriate scalings of the spin operators. For instance: $X_{j}^{(s)}\rightarrow\sqrt{\frac{s+1}{3s}}X_{j}^{(s)}$ (similar for $Y_{j}^{(s)}$ and $Z_{j}^{(s)}$) for the first case Xu:2019lhc ; Bhattacharjee:2022vlt and $X_{j}^{(s)}\rightarrow\sqrt{\frac{s(s+1)}{3}}X_{j}^{(s)}$ (similar for $Y_{j}^{(s)}$ and $Z_{j}^{(s)}$) for the second case Yin:2020oze . These scalings effectively normalize the eigenvalues to align with the conventions outlined in Table 1. Consequently, regardless of the specific construction chosen, the resulting operator statistics remain consistent and yield equivalent results.
6.2 Operator statistics at different times
Figure 6 illustrates the eigenvalue statistics of the time-evolved $s=1/2$ operator across three distinct phases of the system: the ergodic phase $\gamma=0$, the fractal phase $\gamma=1.5$, and the localized phase $\gamma=5$. These specific values of $\gamma$ are selected to ensure consistency with the eigenvalue statistics derived for the respective dimensions of the Hamiltonian; see Fig. 2.
At initial time $t=0$, the operator statistics are independent of the Hamiltonian and thus follow the multinomial distribution, depending on their number of distinct eigenvalues (see Table 1). In the ergodic phase, the eigenvalue statistics of the operator exhibit convergence toward the arcsine distribution at the chosen timestamp, as predicted by free probability theory Camargo:2025zxr . This behavior reflects the expected random matrix-like dynamics within the ergodic regime. In contrast, the eigenvalue distributions in the fractal and localized phases deviate significantly from the arcsine distribution at the same timestamp. These deviations highlight the role of different nature of the quantum dynamics in the fractal and localized regimes, and we can speculate that quantum dynamics in these regimes of the Hamiltonian do not drive the operators (here $S_{1}(0)$ and $S_{L}(0)$) towards freeness (see also the discussion on the late time behavior of these statistics below).
Figure 7: The statistics of the eigenvalues of the time-evolved $s=1$ operator at the initial time $t=0$ (top, left) and late times $t=50$ in the ergodic (top, right), fractal (bottom, left), and localized (bottom, right) regimes. We choose $L=6$ (i.e., dimension of the Hamiltonian $N=(2s+1)^{L}=729$) and average over $1000$ independent realizations. The solid blue lines describe the distribution predicted by the free probability Eq. (14), i.e., Fig. 1 (b).
Figure 8: The statistics of the eigenvalues of the time-evolved $s=3/2$ operator at the initial time $t=0$ (top, left) and late times $t=50$ in the ergodic (top, right), fractal (bottom, left), and localized (bottom, right) regimes. We choose $L=5$ (i.e., dimension of the Hamiltonian $N=(2s+1)^{L}=1024$) and average over $500$ independent realizations of the Hamiltonian. The solid blue lines describe the distribution predicted by the free probability, i.e., Fig. 1 (c).
The same calculation is extended to spin $s=1$ and spin $s=3/2$ operators, as given in (42)-(43). The resulting operator statistics are illustrated in Fig. 7 and Fig. 8, corresponding to $s=1$ and $s=3/2$ respectively. The analytical expression for the distribution predicted by the free probability theory for $s=1$ is given in eq. (14), derived previously in Camargo:2025zxr , while for $s=3/2$ the exact analytical expression of the free probability prediction is not known. The numerical findings for $s=1$ and $s=3/2$ remain consistent with the conclusions derived for the $s=1/2$.
A natural question one can now ask is the following: do the eigenvalue statistics in the fractal and localized phases eventually approach the free probability prediction at sufficiently late times? [^13] The specific shape of the distribution might hold its own intrinsic interest; however, our primary focus is on whether it converges to the free probability distribution. We thank Kohei Kawabata and Hal Tasaki for discussions on this point. The results presented in Fig. 9 suggest that they do not. At very late times $t\sim O(10^{4})$, for all the spins, the operator eigenvalue distribution in the fractal phase stabilizes into a distinct form, which remains fundamentally different from the free probability prediction. However, there are several spikes observed, exactly at the same point, keeping the memory of the initial distribution at $t=0$. This indicates that the fractal phase sustains its intermediate, non-ergodic characteristics even at prolonged timescales.
In the localized phase, the eigenvalue statistics display an even more striking behavior. Specifically, the operator statistics remain almost unchanged from their initial configuration $t=0$, signaling a complete absence of dynamical evolution. This phenomenon aligns with the hallmark of localization, where the degrees of freedom of the system are effectively frozen, preventing significant operator spreading.
Figure 9: The statistics of the eigenvalues of the time-evolved $s=1/2$ (top panel), $s=1$ (middle panel), and $s=3/2$ (bottom panel) operator at the very late times $t\sim O(10^{4})$ in the fractal (left panels) and the localized (right panels) phases. The solid lines in all panels describe the distribution predicted by the free probability. The system parameters are the same as Fig. 6 – Fig. 8. In the fractal regime, the operator statistics converge to a specific distribution that deviates from the free probability prediction. Distinct spikes are observed at the same points as those of at $t=0$, indicating that the time-evolved operator retain some ‘memory’ of the initial distribution. Conversely, in the localized regime, the operator statistics remain markedly different from the results predicted by free probability, indicating the persistence of strong localization effects.
6.3 Analysis of the free time
In the preceding sections, we demonstrated that at late times, in the ergodic phase, the RP model Hamiltonian exhibits operator statistics consistent with predictions from free probability theory. However, the associated timescale predictions have thus far remained predominantly qualitative rather than quantitative. In this section, we aim to bridge this gap by undertaking a rigorous quantification of the timescales. Our approach incorporates a range of methodologies, including the use of distance metrics and advanced statistical techniques like hypothesis testing.
We begin with a two-distance measure, which computes the magnitude of the difference between the histogram data points and the corresponding analytical values. First consider the following function Camargo:2025zxr ,
$\chi^{2}(t)=\sum_{i\neq\mathrm{bdy}}\bigg{(}\frac{\rho(\lambda_{i},t)-\rho_{\mathrm{free}}(\lambda_{i})}{\rho_{\mathrm{free}}(\lambda_{i})}\bigg{)}^{2}\,,$ (52)
where $\rho(\lambda_{i},t)$ is the operator density at time $t$ and $\rho_{\mathrm{free}}(\lambda_{i})$ denotes the eigenvalue density obtained from the free probability prediction. Here $\lambda_{i}$ is the $i$-th bin interval, which sums over the total number of full intervals of the operator statistics. It is important to note that this quantity is slightly different from other well-known Goodness-of-Fit tests corder2014nonparametric ; deshpande2017nonparametric , such as Pearson’s $\chi^{2}$ test, by its tailored focus on operator density deviations.
Another well-known information theoretic distance measure is the Kullback–Leibler (KL) divergence amari2000methods , also known as the relative entropy in the context when $\rho(\lambda)$ is the density matrix of some quantum mechanical system. Since we want to quantify the distance between a numerically obtained distribution and the distribution predicted by the free probability, here we define the KL divergence as follows
$\mathcal{D}_{\mathrm{KL}}(t)\equiv\mathcal{D}_{\mathrm{KL}}(\rho(\lambda_{i},t)||\rho_{\mathrm{free}}(\lambda_{i})):=\sum_{i\neq\mathrm{bdy}}\rho(\lambda_{i},t)\bigg{|}\log[\frac{\rho(\lambda_{i},t)}{\rho_{\mathrm{free}}(\lambda_{i})}\bigg{]}\bigg{|}\,.$ (53)
In this definition, we have taken the absolute value of the term involving the logarithm to ensure that only the relative magnitude ratio of the two corresponding points of the histograms is important. However, note that in the usual definition of KL divergence, only the ratio $\rho(\lambda_{i},t)/\rho_{\mathrm{free}}(\lambda_{i})$ appears, not its absolute value. At earlier times when $\rho(\lambda_{i},t)=0$ for some particular $\lambda_{i}$, the above expression needs to be taken as the limit case such that $\lim_{x\rightarrow 0}x\log x=0$.
Figure 10: The behavior of the the $\chi^{2}$ function and the KL divergence with time for $\gamma=0$ (top, left), $\gamma=0.3$ (top, right), $\gamma=0.5$ (bottom, left), and $\gamma=0.8$ (bottom, right). In each panel, the data points for $\chi^{2}$ lie above the data points for the KL divergence. We choose $s=1/2$, $L=8$, and take averages over $500$ realizations.
Figure 10 illustrates the temporal evolution of the $\chi^{2}$ function and the KL divergence for four distinct values of $\gamma$. The solid gray lines indicate the associated free times for each method, determined by fitting the following exponentially decaying function:
$\mathcal{F}_{\mathrm{fit}}(t):=\mathcal{B}\,e^{-(t-t_{0})/t_{d}}+\mathcal{B}\,,~{}~{}~{}~{}~{}~{}~{}\mathcal{F}_{\mathrm{fit}}=\chi^{2}_{\mathrm{fit}}~{}~{}\mathrm{or}~{}~{}(\mathcal{D}_{\mathrm{KL}})_{\mathrm{fit}}\,,$ (54)
where $\mathcal{B}$, $t_{0}$, and $t_{d}$ are fitting parameters yet to be determined. As $t\rightarrow\infty$, the function $\mathcal{F}_{\mathrm{fit}}(t)\rightarrow\mathcal{B}$, signifying that $\mathcal{B}$ represents the saturation value at late times.
The free time, denoted as $t_{\mathrm{free}}$, represents the timescale that satisfies the following condition:
$t_{\mathrm{free}}:=t_{0}+n\,t_{d}\,,~{}~{}n=1,2,\cdots,~{}~{}~{}\Rightarrow~{}~{}~{}\mathcal{F}_{\mathrm{fit}}(t_{\mathrm{free}})=\mathcal{B}\left(\frac{1}{e^{n}}+1\right)\,.$ (55)
To estimate $t_{\mathrm{free}}$, we identify the timescale at which the difference between $\mathcal{F}_{\mathrm{fit}}(t)$ and its saturation value $\mathcal{B}$ satisfies:
$|\mathcal{F}_{\mathrm{fit}}(t)-\mathcal{B}|=\frac{\mathcal{B}}{e^{n}}\leq\delta\,,$ (56)
where $\delta$ is a prescribed tolerance. Determining $n$ to fulfill this condition and substituting it into (55), we obtain the free time.
It is important to note that at late times, the saturation value of the KL divergence exceeds that of the $\chi^{2}$ function. [^14] We thank Tanay Pathak for asking this question. However, this does not imply that the KL divergence is inferior to the $\chi^{2}$ function. The distinction arises from the very nature of their definitions, as given in (52) and (53). To illustrate this, consider an instance when the operator statistics align with the predictions of free probability but exhibit small numerical fluctuations around the predicted value. Define $\rho_{\mathrm{free}}(\lambda_{i})\simeq u$ and $\rho(\lambda_{i},t)\simeq u+\epsilon$, where $u$ is constant and $\epsilon\ll u$ represents a minor fluctuation relative to $u$. Using these, the respective distance measures evaluate as:
$\chi^{2}$ $\simeq\bigg{(}\frac{(u+\epsilon)-u}{u}\bigg{)}^{2}=\frac{\epsilon^{2}}{u^{2}}\,,$
$\mathcal{D}_{\mathrm{KL}}$ $\simeq(u+\epsilon)\bigg{|}\log[\frac{u+\epsilon}{u}\bigg{]}\bigg{|}=\epsilon+\frac{\epsilon^{2}}{u}\,.$ (57)
Given that $\epsilon\ll u$, it follows that $\chi^{2}\simeq O(\epsilon^{2})$ at leading order, while $\mathcal{D}_{\mathrm{KL}}\simeq O(\epsilon)$. This establishes that by definition, the saturation value of $\chi^{2}$ will be lower than that of the KL divergence, justifying the results in Fig. 10.
Figure 11 (a) shows the behavior of the free time with respect to $\gamma$ (brown for $\chi^{2}$ function and red for KL divergence). We chose $n=2$ in (55) such that the error remains within $10\%$ of the saturation value. As evident, the time increases as $\gamma$ increases, i.e., the operator statistics lead to the free probability prediction at later times as the system moves away from the ergodic regime. The dependence of free time on $L$ (the system size being $2^{L}$) from $L=6$ to $L=12$ is shown in Fig. 11 (b) for both distance measures. The error estimates are computed by the standard error estimates in Mathematica. The errors for the $\chi^{2}$ function and KL divergence are of $O(10^{-3})$ and $O(10^{-2})$ respectively.
Figure 11: (a) The free time (or the arcsine time for spin $1/2$) with the deformation $\gamma$, computed from the $\chi^{2}$ function (brown), KL divergence (red), and the Kolmogorov-Smirnov (KS) test (blue). (b) The finite-size scaling of free time for a fixed $\gamma=0$. The corresponding realizations are taken as $L=6\,(5000)$, $L=7\,(3000)$, $L=8\,(1000)$, $L=9\,(500)$, $L=10\,(200)$, $L=11\,(100)$, and $L=12\,(50)$. The dimension of the Hilbert space is $2^{L}$. The errors are computed by the standard error estimates in Mathematica.
The final method we consider is the Kolmogorov-Smirnov (KS) test kolmogorov1933 ; smirnov1939estimation ; smirnov1948 , a non-parametric statistical test used to determine whether two one-dimensional probability distributions are identical. It is commonly employed to assess whether empirical data could plausibly originate from a given reference distribution. In the language of hypothesis testing, the assumption that both distributions are equal defines the null hypothesis, denoted by $H_{0}$, while the alternative hypothesis, $H_{1}$, corresponds to the case in which they differ corder2014nonparametric .
Let us consider an observed sample ${\lambda_{1},\lambda_{2},\cdots,\lambda_{N}}$ consisting of $N$ data points. We aim to determine whether this sample follows a specific reference probability density function, denoted by $\text{PDF}_{\text{ref}}(\lambda)$. The KS test compares the empirical cumulative distribution function (CDF) of the sample with the CDF of the reference distribution. The latter is computed as
[…]
The empirical CDF is defined as the proportion of sample points less than or equal to a given value $\lambda$, and can be written as
$\text{CDF}_{N}(\lambda)=\frac{1}{N}\sum_{i=1}^{N}\theta(\lambda-\lambda_{i})\,,$ (59)
where $\theta$ is the Heaviside step function.
Figure 12: The temporal evolution of $p$-value for the deformation $\gamma$ from the KS test. The red dashed line indicates the value of $p=0.05$, above which the null hypothesis is accepted, and the empirical distribution of the operator follows the free probability prediction. The grey dashed line indicates the free time, where the curves intersect the red dashed line.
The KS statistic is then defined as the maximum absolute difference between the empirical and reference CDFs across all $\lambda$ values corder2014nonparametric :
$D_{N}=\sup_{\lambda}|\text{CDF}_{N}(\lambda)-\text{CDF}_{\text{ref}}(\lambda)|\,.$ (60)
Under the assumption that the null hypothesis is true, and in the limit of large $N$, the rescaled statistic $\sqrt{N}D_{N}$ converges in distribution to a random variable $x$ governed by the Kolmogorov distribution $\text{PDF}_{\mathrm{K}}$:
$\lim_{N\rightarrow\infty}\sqrt{N}D_{N}\rightarrow x\sim\text{PDF}_{\mathrm{K}}(x)\,.$ (61)
The Kolmogorov probability density function is given by
$\text{PDF}_{\mathrm{K}}(x)=\frac{d}{dx}\text{CDF}_{\mathrm{K}}(x)=8x\sum_{n=1}^{\infty}(-1)^{n-1}n^{2}e^{-2n^{2}x^{2}}\,.$ (62)
The associated cumulative distribution function is given by KStest
[…]
The procedure of the KS test involves computing the statistic $x=\sqrt{N}D_{N}$ and evaluating whether it is significantly large. This is done by evaluating the associated $p$-value $=1-\text{CDF}_{\text{K}}(x)$, which is the probability of finding a larger value for the parameter $x=\sqrt{N}D_{N}$ given that the null hypothesis is true (the two distributions are the same). If the $p$-value is smaller than 0.05, one concludes that the parameter is too large and the null hypothesis is false, namely, the two distributions are not the same. Otherwise, if the $p$-value is larger than 0.05, one concludes that the null hypothesis is true and the two distributions are the same. [^15] Given two distributions, the KS test can be easily implemented in Mathematica by using the command KolmogorovSmirnovTest. Figure 13 illustrates the Kolmogorov PDF and the $p$-value associated with a particular value of $x=\sqrt{N}D_{N}$.
Figure 13: Kolmogorov probability density function (PDF). The shaded area illustrates the $p$-value associated with a particular value of $x=\sqrt{N}D_{N}$. In hypothesis testing, a $p$-value below a chosen significance level – commonly 0.05 – indicates that the null hypothesis (that the two distributions are the same) can be rejected with high confidence. For the value $\sqrt{N}D_{N}=1.1$ shown by the red dashed line, the $p$-value equals 0.178. In this case, the two distributions would be considered the same.
We apply the KS test to assess whether the empirical distributions for the eigenvalues of $A(0)+A(t)$ follow the prediction from free probability. In Fig. 12, we show the $p$-value of the test as a function of time for different values of $\gamma$. The figure shows that at early times the $p$-value takes values smaller than 0.05, indicating that the empirical distribution and the free probability prediction are not the same. However, as time increases, the $p$-value increases, taking values larger than 0.05, indicating that the free probability prediction has emerged. The free time is computed as the time at which the $p$-value crosses the value 0.05.161616We also checked with other tests such as Pearson’s $\chi^{2}$ test and confirm the strikingly similar behavior of free time, as in Fig. 11.
7 Conclusion and outlook
Within the framework of free probability, we revisit the calculation of the DOS for the Rosenzweig–Porter (RP) random matrix ensemble and study the spectral statistics of operators, highlighting the approximate emergence of asymptotic freeness in this model.
Since the RP model is defined as the sum of two independent random matrices, the two terms are free in the sense of free probability in the limit when the dimension of the matrices is sufficiently large. This enables the use of free probability techniques to compute the DOS. Using a general approach based on subordination functions, we derive an explicit expression for the DOS of the RP model, reproducing previous results obtained via the replica technique Venturelli:2022hka . Our analysis focuses on the fractal phase, where we express the DOS as perturbative corrections to the Gaussian profile characteristic of the localized phase. As illustrated in Fig. 3, the DOS evolves smoothly from a Wigner semicircle to a Gaussian distribution as the parameter $\gamma$ increases. In particular, during the transition from the ergodic to the fractal phase, the DOS resembles a perturbed Wigner semicircle. It would be interesting to extend the methodology developed here to compute the DOS in this intermediate regime.
It has been recently proposed that chaotic dynamics drives the components of a sum of two operators, one of which is time-evolved, toward asymptotic freeness with respect to non-evolved ones Camargo:2025zxr . However, the precise mechanisms underlying the emergence of freeness remain unclear. For instance, is it sufficient for the spectral statistics of the Hamiltonian to follow random matrix theory? What role do the eigenvectors play in this process? To address these questions, we studied the emergence of free probability predictions for the spectrum of operators of the form $S_{1}+S_{L}(t)$ in the RP model, an ensemble of random matrix Hamiltonians where the rotational invariance of the classical Gaussian ensembles is broken explicitly by introducing a parameter, and depending on the value of this parameter the energy eigenvectors show different types of localization properties. Our analysis indicates that such behavior appears only in the ergodic phase; in contrast, free probability fails to accurately describe the operator spectrum in the fractal and localized phases, as expected.
Interestingly, in the fractal phase, the system exhibits partial signatures of freeness, yet retains memory of the spectrum at $t=0$. This suggests that the emergence of freeness is not solely governed by spectral statistics, but also by the structure of the eigenvectors. While the fractal phase still displays level repulsion—as reflected in the average $r$-parameter—the corresponding eigenvectors are not fully delocalized. This points to a natural conclusion: the presence of level repulsion in the Hamiltonian alone is not sufficient for the emergence of freeness in operator statistics; extended, non-localized eigenvectors are also essential. It would be interesting to explore this direction further and test its validity in other models that exhibit fractal or multifractal phases Khaymovich:2021tkj .
We have utilized various measures to accurately characterize the timescale at which free probability predictions emerge, marking the onset of free time. These methods include distance metrics like the $\chi^{2}$ function and KL divergence, alongside statistical techniques such as hypothesis testing. Notably, all methods yield consistent and reliable results. Additionally, the agreement between the KS test (based on the CDF) and both the KL divergence and $\chi^{2}$ test (based on the PDF) underscores the robustness of PDF-based methods in our analysis. This consistency confirms that the choice of histogram bin size does not affect our findings within the estimated error ranges, a conclusion supported by the considerable number of Hamiltonian ensembles in the operator statistics evaluation.
Several compelling directions emerge for future research. Notably, the sharp transition observed in the RP model manifests in the infinite-size limit, precisely the regime where predictions from free probability theory become exact. This alignment makes it particularly intriguing to attempt to identify the transition point within a free probability framework. A promising approach might involve fixing a sufficiently large time and analyzing the variation of the KL divergence with respect to $\gamma$, which measures the distance between the empirical distribution and the free probability distribution. Such an analysis could potentially reveal whether this distance provides a clear demarcation for the distinct transitions within the model. We expect to report on this in future work.
In the present study, our focus has been restricted to the statistical properties of eigenvalues. However, the role of eigenvectors is equally crucial, especially in the context of thermalization. A natural extension would be to investigate the localization and delocalization behavior of the eigenvectors of the operator $S_{1}+S_{L}(t)$, which could shed light on the mechanisms underlying the ETH. Such an investigation would be especially illuminating when approached through the framework of free cumulants, potentially revealing interesting connections with recent developments in the field Pappalardi:2022aaz ; Pappalardi:2023nsj ; Fritzsch:2024qjn ; Wang:2025mzl ; Vallini:2024bwp ; Fava:2023pac , with our results on operator statistics. Especially, it would be an excellent direction to compare the free time studied here with the timescale for the onset of freeness introduced in Ref. Vallini:2024bwp , generalizing to systems hosting multifractal states Khaymovich:2021tkj . Pursuing this direction would require the consideration of finite-temperature states, thereby necessitating a generalization of the current statistical framework. Addressing these questions will be the subject of future investigations.
Acknowledgements
We thank Yichao Fu, Kohei Kawabata, Tokiro Numasawa, Tanay Pathak, Hal Tasaki, and Masafumi Udagawa for fruitful discussions. We also thank Jonah Kudler-Flam for useful comments on the draft. P.N. acknowledges the hospitality of Gwangju Institute of Science and Technology (GIST), Gakushuin University and the Institute of Solid State Physics (ISSP), University of Tokyo, during the final stages of the work. This work was supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Science, ICT & Future Planning (NRF-2021R1A2C1006791, NRF-2020R1I1A2054376), the framework of international cooperation program managed by the NRF of Korea (RS-2023-NR119907, RS-2024-00440725), the Creation of the Quantum Information Science R&D Ecosystem (Grant No. 2022M3H3A106307411) through the NRF funded by the Korean government (Ministry of Science and ICT), the Gwangju Institute of Science and Technology (GIST) research fund (Future leading Specialized Resarch Project, 2025) and the Al-based GIST Research Scientist Project grant funded by the GIST in 2025. V.J. and H.C. are supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (NRF-2022R1I1A1A01070589, and RS-2023-00248186). The work of P.N. is supported by the JSPS Grant-in-Aid for Transformative Research Areas (A) “Extreme Universe” No. 21H05190.
Appendix A Comparison of Haar random unitaries and exponential of Gaussian ensembles
In this appendix, we compare the asymptotic freeness governed by different $N\times N$ unitary ensembles. The unitary ensembles considered are listed below:
$U_{\mathrm{CXE}}\,,~{}~{}~{}U_{\mathrm{GXE}}=e^{-iH_{\mathrm{GXE}}t}\,,~{}~{}~{}\mathrm{X}\in\{\mathrm{O},\mathrm{U}\}\,,$ (64)
with $t\sim O(1)$. This list includes the corresponding orthogonal and unitary ensembles. The conclusions for the symplectic ensemble can also be generalized straightforwardly. Among the unitaries in (64), $U_{\mathrm{CXE}}$ with $\mathrm{X}\in\{\mathrm{O},\mathrm{U}\}$ are the Haar random, while $U_{\mathrm{GXE}}=e^{-iH_{\mathrm{GXE}}t}$ with $t\sim O(1)$ are not. Following Cotler:2017jue , we refer to the latter as an ensemble of Gaussian Hamiltonians. This distinction can be easily verified by examining the spectrum of the phase of the eigenvalues of the corresponding unitaries. More specifically, for all cases, the eigenvalues are of the form $\lambda_{k}=re^{i\theta_{k}}$, where $\theta_{k}=\mathrm{Arg}(\lambda_{k})\in[-\pi,\pi]$ and $k=1,\cdots,N$. Therefore, we consider the spacing distribution of $\theta_{k}$, defined as $\xi_{j}=\frac{N}{2\pi}(\theta_{j+1}-\theta_{j})$ with $j=1,\cdots,N-1$, and the $r$-ratio, defined in (18) for all unitaries (64). While the DOS $\rho(\theta)=1/(2\pi)$ is uniform, the unitaries $U_{\mathrm{CXE}}$ exhibit level repulsion while $U_{\mathrm{GXE}}$ show behavior typical of an uncorrelated ensemble. The nearest-level spacing distribution $\xi_{k}$ and the $r$-ratio distribution are illustrated in Fig. 14 and Fig. 15 respectively. The analytic results are matched with (17) and (19) respectively.
Figure 14: Nearest-level spacing distribution for the unitaries $U_{\mathrm{COE}}$ (left), $U_{\mathrm{CUE}}$ (middle) and $U_{\mathrm{GXE}}$ (right), with $\mathrm{X}\in\{\mathrm{O},\mathrm{U}\}$ (both produce identical plots) for $s=1/2$. The red lines are the analytic results for $\upbeta=1,2$ and 00 respectively (from left to right). We choose $L=8$ and average over $5000$ samples. For $U_{\mathrm{GXE}}$, we choose $t=50$.
Figure 15: Histogram of the $r$-ratio for the unitaries $U_{\mathrm{COE}}$ (left), $U_{\mathrm{CUE}}$ (middle) and $U_{\mathrm{GXE}}$ (right), with $\mathrm{X}\in\{\mathrm{O},\mathrm{U}\}$ (both produce identical plots) for $s=1/2$. The red lines are the analytic results for $\upbeta=1,2$ and 00 respectively (from left to right). We choose $L=8$ and average over $5000$ samples. For $U_{\mathrm{GXE}}$, we choose $t=50$.
Next, we perform the operator evolution with all unitaries in (64). Interestingly, we find that all of them drive the operators to asymptotic freeness. Figure 16 shows the comparison between the Haar random ensembles and the exponential of the Gaussian ensembles. For simplicity, we only show the results for $s=1/2$, but $s=1$ and $s=3/2$ can be obtained straightforwardly. This illustrates that the asymptotic freeness is a more generic property and need not be tied to the Haar randomness of the unitary operator in the Hamiltonian evolution. It will be worthwhile to examine what properties of the unitary are responsible for freeness.
Figure 16: The operator statistics governed by the unitaries $U_{\mathrm{COE}}$ (top) and $U_{\mathrm{GOE}}$ (bottom) for $s=1/2$. We choose $L=8$ and average over $5000$ samples. The blue solid lines indicate the free probability result, i.e., the arcsine distribution (12). Both unitaries, although they exhibit different level statistics (left and right in Fig. 15), drive the operators $S_{1}$ and $S_{L}$ to asymptotically free, so the plots become indistinguishable from each other.
Appendix B Free additive convolution of spin-$s$ operators
In this appendix, we review the derivation of the free probability prediction for the spectrum of the sum of two free spin-$s$ operators, whose spectral distributions take the form:
$\rho_{s}(\lambda)=\frac{1}{2s+1}\sum_{j=-s}^{s}\delta(\lambda-s)\,.$ (65)
The corresponding Cauchy transform is given by
$G_{s}(z)=\int\frac{\rho_{s}(\lambda)}{z-\lambda}=\frac{1}{2s+1}\sum_{j=-s}^{s}\frac{1}{z-s}\,.$ (66)
The associated R-transform, defined as $R_{s}(z)=G^{-1}_{s}(z)-1/z$, is obtained by solving an algebraic equation of order $s$. This poses a practical challenge for large values of $s$: for example, computing the R-transform for spin-$3/2$ operators already involves solving a quartic equation, with even higher-degree equations required for larger spin values. It thus becomes evident that one must resort to numerical methods in such cases.
A particularly effective numerical technique for evaluating the free additive convolution of random variables with arbitrary measures is the subordination function method. We refer to Sec. 5 of speicher2019lecturenotesfreeprobability for a thorough discussion of this method, including detailed derivations of all formulas used here.
Suppose we are given two free random variables $a$ and $b$, with Cauchy transforms $G_{a}(z)$ and $G_{b}(z)$, respectively. It can be shown that the Cauchy transform of their sum $\rho_{a}\boxplus\rho_{b}$ is given by speicher2019lecturenotesfreeprobability
$G_{\rho_{a}\boxplus\rho_{b}}(z)=G_{a}(\omega_{a}(z))\,,$ (67)
where the subordination function $\omega_{a}(z)$ satisfies the fixed-point equation for $\omega_{a}(z)\in\mathbb{C}^{+}$ (the upper half-plane):
$\omega_{a}(z)=z+H_{b}\left[z+H_{a}(\omega_{a}(z))\right]\,,$ (68)
with
$H_{a}(z)=\frac{1}{G_{a}(z)}-z\,,~{}~{}~{}~{}H_{b}(z)=\frac{1}{G_{b}(z)}-z\,,$ (69)
being auxiliary functions. Finally, the spectral distribution of $a+b$ is obtained via the inverse Stieltjes transform of (67):
$\rho_{a+b}(\lambda)=-\lim_{\epsilon\rightarrow 0}\frac{1}{\pi}\text{Im}\,G_{\rho_{a}\boxplus\rho_{b}}(\lambda+i\epsilon)\,.$ (70)
Now we illustrate the application of this method by using it to compute the spectral density of the sum of two spin operators with $s=10$. This value is chosen to demonstrate that the method remains applicable for relatively large $s$, although the resulting expressions become increasingly complicated as $s$ increases.
Figure 17: Spectral distribution of the sum of two spin-10 operators, $a=S_{1}$ and $b=U^{\dagger}S_{1}U$, where $U$ is a Haar-random unitary and $S_{1}$ is defined in Eq. (43) with $s=10$ and $L=2$. The solid line shows the free probability prediction, compared with histogram data obtained from the direct diagonalization of $a+b$ over $10^{4}$ realizations.
Example: Sum of two free spin-10 operators:
Consider the sum of two free spin-10 operators, $a=S_{1}$ and $b=U^{\dagger}S_{1}U$, where $U$ is a Haar-random unitary and $S_{1}$ is defined in Eq.(43) for $s=10$ and $L=2$. Both operators share the same spectral distribution and Cauchy transform, given by (65) and (66) for $s=10$. The corresponding auxiliary functions are:
$H_{a}(z)=H_{b}(z)=\frac{P(z)}{Q(z)}\,.$ (71)
where
$P(z)$ $=2z\Big{(}\big{(}-41309657025z^{6}+911087234376z^{4}-11156038265680z^{2}\big{)}z^{4}$
$\quad-183668715648000z^{2}+-11\big{(}35z^{6}-11172z^{4}+1443810z^{2}-97774684\big{)}z^{12}$
$\quad+69514526366208z^{4}+131681894400000\Big{)}\,,$ (72)
and
$Q(z)$ $=21z^{20}-7315z^{18}+1044582z^{16}-79409550z^{14}+3495444953z^{12}$
$\quad-90881245455z^{10}+1366630851564z^{8}-11156038265680z^{6}$
$\quad+43446578978880z^{4}-61222905216000z^{2}+13168189440000\,.$ (73)
After numerically solving (68) for $H_{a}$ and $H_{b}$ given by (71), and using (70), we obtain the spectral density $\rho_{a+b}(\lambda)$ numerically. The result is shown in Fig. 17 and compared to the numerical data obtained by direct diagonalization of $a+b$.